Total number of ways of factorising the number of 676 in two factors,that are prime to each other is equal to..........

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi,

on factorising 676 = 2^2 * 13^2

so, we get to prime numbers 2 & 13

 for gettig two prime factors, we should take care of that there happens no mixing of these prime no.

you can treat 2^2 as A & 13 ^2 as B 

then ,  676 = A * B

its clear that only two cases can happen,

1) A =1 , B = 676

2) A = 2^2  , B =  13 ^2

ans ... 2 no. of ways are possible